Circular No. 8076

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

IGR J16320–4751 J. A. Tomsick and R. Lingenfelter, University of California at San

Diego; R. Walter, INTEGRAL Science Data Centre; J. Rodriguez and A. Goldwurm, Commissariat à l’Energie Atomique (CEA), Saclay; S. Corbel, Université Paris 7 and CEA, Saclay; and P. Kaaret, Harvard-Smithsonian Center for Astrophysics, report the detection of a variable hard x-ray source, IGR J16320–4751, with the IBIS instrument’s ISGRI detector onboard the INTEGRAL satellite. The source was detected at α = 16^{h}32^{m}.0, δ = −47^{o}51′ (equinox 2000.0; uncertainty 2′) during an observation of 4U 1630–47 on Feb. 1.4 UT. The 15–40-keV source flux varies from < 10 to 50 mCrab on a time scale of 1000 s, and the source was detected above 60 keV. The position is consistent with that of the source AX J1631.9–4752 (detected during ASCA observations in 1994 and 1997). The ASCA/GIS source count rate varies by a factor of 4 during the 14000-s 1997 observation, and Sugizaki et al. (2001, Ap.J. Suppl. 134, 77) fit the GIS energy spectrum with an absorbed power law with a photon index of 0.2 ± 0.2 and a hydrogen column density of (9 ± 1) × 10^{22} cm^{−2}, finding a mean 0.7–10-keV flux near 5 × 10^{−11} erg cm^{−2} s^{−1}. The hard power-law index may suggest that the source is a high-mass x-ray binary, indicating that it may be possible to identify the source at other wavelengths.

IGR J16318–4848 L. Foschini, Istituto di Astrofisica Spaziale e Fisica Cosmica, Consiglio

Nazionale delle Ricerche, Bologna; J. Rodriguez, Service d’Astrophysique, CEA, Saclay; and R. Walter, INTEGRAL Science Data Centre, Geneva, report on the Digitized Sky Survey (DSS) II, 2MASS, and Midcourse Space Experiment (MSX) archival data analyses of the field of IGR J16318–4848 (IAUC 8063). A clear source is detected in the 2MASS All-Sky Quicklook image archive in the three bands J, H, and K, at α = 16^{h}31^{m}48^{s}.3, δ = −48^{o}49′01′′ (equinox 2000.0; nominal uncertainty ∼ 0′′.2) — consistent with ASCA (IAUC 8070) and XMM positions (IAUC 8072). Very rough magnitude estimates (1-mag uncertainties): J = 10, H = 9, K = 9. A counterpart is also clearly detected in the I band of the second DSS at position end figures 48^{s}.2, 49′00′′. No counterpart is visible in the R band of the second DSS, to R = 21. A source is also visible in the MSX5C catalogue at position end figures 48^{s}.3, 48′57′′ (uncertainty 1′′.5), consistent with that of 2MASS. The flux density at 8 μm is 0.4557 Jy. The above data confirm that the source is heavily absorbed.

2003 February 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green